### 30 Mar 2026 v2

# Lee–Yang Zeros and Pseudocritical Drift in J-Q Néel–VBS Transitions

Abstract
Square-lattice J-Q models provide a sign-problem-free setting for probing the quantum phase transition between Néel antiferromagnet and columnar valence-bond solid. We analyze this transition through the scaling of Lee–Yang zeros, computed within stochastic series expansion quantum Monte Carlo by reweighting configurations sampled near criticality in the presence of complex source fields. Benchmark studies of the dimerized Heisenberg model and the checkerboard J-Q model validate the method, yielding stable $\mathrm{O}(3)$ critical scaling in the former and clear spacetime-volume scaling in the latter, as expected for a first-order transition. Applying the same analysis to the J-Q models, we find a pronounced and systematic drift of the leading-zero scaling with increasing system size, consistent with an extended pseudocritical regime. The Lee–Yang scaling implies an effective scaling dimension $\Delta_{\phi}$ of the $\mathrm{SO}(5)$ order-parameter field that decreases with size and is consistent with vanishing in the thermodynamic limit. Such behavior lies below the scalar unitarity bound of any unitary relativistic conformal field theory in 2+12{+}1 dimensions and enforces inverse spacetime-volume scaling of the zeros, the hallmark of a first-order transition. These results support a weakly first-order interpretation of the Néel–VBS transition and establish the finite-size Lee–Yang zeros as a sensitive, symmetry-resolved diagnostic of pseudocriticality and transition order in the J-Q family.
Introduction. Quantum magnets provide a controlled setting for studying quantum phase transitions beyond the standard Landau–Ginzburg–Wilson framework based on a single local order parameter [65, 12, 75, 74, 76]. A central example is the square-lattice transition between Néel antiferromagnet (AFM) and a columnar valence-bond solid (VBS), realized in sign-problem-free J-Q models and accessible to large-scale quantum Monte Carlo simulations [70, 57, 53, 72, 78, 60, 68, 84]. Within conventional Landau theory, a direct continuous transition between these phases is not generically expected without additional fine tuning, since the Néel and VBS phases break distinct symmetries [12, 65, 75, 74]. The deconfined quantum critical point (DQCP) scenario instead posits a continuous transition governed by fractionalized spinons coupled to an emergent gauge field, with the Néel order parameter arising as a spinon bilinear and the VBS order associated with monopole operators [75, 74, 43, 36, 76, 63, 77, 55, 83, 94, 15, 107, 14, 104, 47, 79, 86, 87, 13]. Complementary to continuum field-theory approaches, Lee–Yang theory characterizes phase transitions through the analytic structure of the partition function $Z$ in the complex plane, where thermodynamic singularities are tied to the accumulation of zeros [103, 42, 32, 4, 85, 49, 31, 16, 92, 45, 46, 39, 90, 89, 91, 29, 3, 1, 22, 100, 48, 27, 2]. Importantly, Lee–Yang theory is naturally formulated at finite size and is therefore well suited to numerical simulations. For finite linear system size $L$ and inverse temperature $\beta$, Lee–Yang zeros are discrete and remain off the real axis, but their scaling with $L$ provides a sharp, nonperturbative diagnostic of transition order and critical behavior [23, 32, 34, 38, 6, 17, 92, 51, 50, 52, 58]. Lee–Yang zeros therefore provide a natural probe of the debated Néel–VBS criticality in J-Q models [72, 78, 60, 105, 56, 73, 35, 40, 84].
Early quantum Monte Carlo studies of square-lattice $\mathrm{SU}(2)$ J-Q models found nearly size-independent crossings of Binder cumulants and scaled correlation length, consistent with a putative continuous Néel–VBS transition [70, 57, 53, 72]. At the same time, conventional finite-size diagnostics in these models remain difficult to interpret [72, 59, 78, 56, 68, 84]. Crossing points can look almost size independent on moderate lattices, yet continue to drift as larger sizes are included, while exponents from standard finite-size fits depend noticeably on the fitting window, indicating that the asymptotic regime is not cleanly reached. Joint AFM–VBS order-parameter distributions also exhibit approximate emergent symmetry, which is suggestive but does not by itself distinguish true criticality from a long-lived pseudocritical regime [59, 60, 93, 76]. Conventional real-axis observables therefore provide important evidence, but do not directly isolate the finite-size singularity scale or decisively determine the character of the transition. Distinguishing a stable unitary fixed point from an anomalously weak first-order transition thus remains a central challenge for understanding Néel–VBS criticality in the J-Q family [36, 59, 78, 56, 76, 96, 108].
In this Letter we study the Néel–VBS transition through Lee–Yang zeros computed within stochastic series expansion (SSE) quantum Monte Carlo [69, 82, 21, 71, 101, 102, 16, 92]. By evaluating complex generating functions in symmetry-resolved channels, we track the finite-size flow of the low-lying Lee–Yang zeros, which directly measures the singularity scale and distinguishes critical behavior from first-order coexistence [32, 62, 5, 9, 6, 7]. Using amplitude-free two-size estimators constructed from low-lying zeros [32, 34, 38, 17, 92], we analyze benchmark models together with the uniform J-Q3 and J-Q2 systems. While the benchmarks recover their expected behavior, J-Q2 and J-Q3 exhibit pronounced finite-size drift, indicating an extended pseudocritical regime. These results support a weakly first-order interpretation of the Néel–VBS transition and highlight the advantage of Lee–Yang zeros as a direct finite-size diagnostic of slow drift [59, 78, 28, 56, 76, 84].
Figure 1: Schematic illustrations of the lattice Hamiltonians studied in this work. (a) Columnar dimerized Heisenberg model with alternating strong and weak nearest-neighbor bonds. (b) Checkerboard J-Q (CBJQ) model with $Q$ interactions on a checkerboard plaquette pattern. (c) Uniform J-Q2 model with $Q_{2}$ interactions on every plaquette. (d) Uniform J-Q3 model with $Q_{3}$ interactions on $2\times 3$ and $3\times 2$ rectangles.
Methods and Models. For finite $L$ and inverse temperature $\beta$, a symmetry-preserving anchor Hamiltonian $H_{0}$ defines an ensemble that respects all microscopic symmetries, so phases and criticality are inferred from the $L$ dependence of symmetry-resolved fluctuations [41, 12, 65, 36]. We probe a channel specified by operator support (site, bond, or plaquette) and translation sector $\mathbf{Q}$, defined from a local density $O(x)$ at position $\mathbf{R}_{x}$ by the momentum projection $O_{\mathbf{Q}}\equiv\sum_{x}e^{\mathrm{i}\mathbf{Q}\cdot\mathbf{R}_{x}}O(x)$, which is a translation eigen-operator and does not mix with inequivalent $\mathbf{Q}$ sectors in correlators [88, 41, 12, 65]. We introduce an imaginary source $\mathrm{i}g_{\mathbf{Q}}$ coupled to $O_{\mathbf{Q}}$, $H(\mathrm{i}g_{\mathbf{Q}})=H_{0}-\mathrm{i}g_{\mathbf{Q}}O_{\mathbf{Q}}$, and work with the normalized generator $G_{L,\beta}(\mathrm{i}g_{\mathbf{Q}})\equiv Z_{L,\beta}(\mathrm{i}g_{\mathbf{Q}})/Z_{L,\beta}(0)$, whose finite-$L$ zeros encode the analytic structure of the complex $Z$ [103, 42, 24, 32, 4] and whose thermodynamic pinching yields free-energy nonanalyticities [103, 42, 24, 32]. [^1] For the formal scaling discussion we adopt the convention $H(\mathrm{i}g_{\mathbf{Q}})=H_{0}-\mathrm{i}g_{\mathbf{Q}}O_{\mathbf{Q}}$. In the explicit SSE implementation the effective sign of the imaginary deformation is channel dependent. These differences amount only to channel-dependent relabelings of the scan coordinate ($g\to-g$ where needed) and therefore do not affect the Lee–Yang zero set or its finite-size scaling. We evaluate $G$ in SSE by reweighting configurations sampled from the real anchor ensemble [16, 18, 95, 19, 20, 97]: for each stored configuration $\mathcal{C}$ we form the ratio $R(\mathrm{i}g_{\mathbf{Q}};\mathcal{C})\equiv W(\mathcal{C};\mathrm{i}g_{\mathbf{Q}})/W(\mathcal{C};0)$, where $W$ is the SSE weight including the source-dependent diagonal matrix elements [69, 82, 71], and compute $G(\mathrm{i}g_{\mathbf{Q}})=\langle R(\mathrm{i}g_{\mathbf{Q}};\mathcal{C})\rangle_{0}$. Lee–Yang zeros are then the values of $\mathrm{i}g_{\mathbf{Q}}$ at which this estimator vanishes [16, 92].
As a continuous-transition example, we study the 2+12{+}1D columnar dimerized Heisenberg model [98, 99], shown in Fig. 1(a),
$H_{\rm dim}=J_{1}\!\!\sum_{\langle ij\rangle\in{\rm weak}}\mathbf{S}_{i}\!\cdot\!\mathbf{S}_{j}+J_{2}\!\!\sum_{\langle ij\rangle\in{\rm strong}}\mathbf{S}_{i}\!\cdot\!\mathbf{S}_{j},$ (1)
which hosts an $\mathrm{O}(3)$ quantum phase transition between a Néel AFM and a quantum-disordered dimer phase [98, 99, 54]. We probe the Néel channel at $\mathbf{Q}=(\pi,\pi)$ through $M^{z}=\sum_{\mathbf{r}}(-1)^{x+y}S^{z}_{\mathbf{r}}$, by adding the imaginary source term $\mathrm{i}hM^{z}$, with $h\in\mathbb{R}$. For each SSE configuration, we record the numbers of diagonal bond operators acting on the two antiparallel local spin configurations $(\uparrow_{i}\downarrow_{j})$ and $(\downarrow_{i}\uparrow_{j})$, denoted $n^{(s)}_{\uparrow\downarrow}$ and $n^{(s)}_{\downarrow\uparrow}$, respectively. [^2] These counts correspond to diagonal SSE vertices in bond sector $s$, with local bond states $(\uparrow_{i}\downarrow_{j})$ or $(\downarrow_{i}\uparrow_{j})$. The bond endpoints $i\in A$ and $j\in B$ on the bipartite lattice are labeled by $s=\mu\epsilon$, where $\mu\in\{x,y\}$ is the bond orientation and $\epsilon=\pm$ denotes the translation-related sublattices, selected by the columnar staggering factor $(-1)^{r_{\mu}}$ [16, 98]. With real anchor denominators $J_{s}$ and $\alpha_{s}=2/(N_{c}J_{s})$ ($N_{c}=4$ on the square lattice) [67, 69], the per-configuration ratio factor is $R_{h}(\mathrm{i}h)=\prod_{s}\bigl(1-\mathrm{i}\alpha_{s}h\bigr)^{n^{(s)}_{\uparrow\downarrow}}\bigl(1+\mathrm{i}\alpha_{s}h\bigr)^{n^{(s)}_{\downarrow\uparrow}}.$
As a controlled first-order benchmark, we use the checkerboard J-Q (CBJQ) model [105, 80, 44], shown in Fig. 1(b),
$H_{\rm CBJQ}=-J\sum_{\langle ij\rangle}P_{ij}-Q\!\!\sum_{\langle ij,kl\rangle\in\square^{\prime}}P_{ij}P_{kl},$ (2)
where $P_{ij}\equiv\tfrac{1}{4}-\mathbf{S}_{i}\!\cdot\!\mathbf{S}_{j}$, and $\square^{\prime}$ denotes one of the two checkerboard plaquette subfamilies, for example the red shaded plaquettes in Fig. 1(b) [105, 80, 36]. This model exhibits a discontinuous transition between a Néel AFM and a plaquette-singlet-solid (PSS) phase [105, 80, 44]. Motivated by the emergent $\mathrm{O}(4)$ symmetry reported at the transition [105, 80], we probe Lee–Yang zeros for both the Néel and PSS order parameters. For the AFM channel, we couple an imaginary source $\mathrm{i}h$ to the staggered magnetization $M^{z}=\sum_{\mathbf{r}}(-1)^{x+y}S^{z}_{\mathbf{r}}$. To detect PSS order, we resolve plaquettes by checkerboard and row parities, $\epsilon_{\rm row}(\square^{\prime})=(-1)^{y(\square^{\prime})}$ [105, 80], and define $O^{\rm PSS}\equiv\sum_{\square^{\prime}}\epsilon_{\rm row}(\square^{\prime})\,\mathcal{O}(\square^{\prime}),$ where $\mathcal{O}(\square^{\prime})$ is the local plaquette density associated with the microscopic interaction. [^3] In the SSE implementation, $\mathcal{O}(\square^{\prime})$ corresponds to the four-spin operator $P_{ij}P_{kl}$ on an active checkerboard plaquette (Eq. (2)). The PSS estimator is obtained by counting $Q$-vertex insertions with row parity $\epsilon_{\rm row}=\pm 1$, denoted $N_{Q+}$ and $N_{Q-}$, which correspond to plaquette-operator insertions selected by the row-staggering factor. In the PSS channel, we introduce an imaginary source term $\mathrm{i}\eta\,O^{\rm PSS}$. For each SSE configuration, the corresponding reweighting factor is determined by the real anchor coupling $Q$ and by the row-parity-resolved numbers of $Q$-vertex insertions, $N_{Q+}$ and $N_{Q-}$, giving $R_{\mathrm{pss}}(\mathrm{i}\eta)=\left(1+\mathrm{i}\frac{\eta}{Q}\right)^{N_{Q+}}\left(1-\mathrm{i}\frac{\eta}{Q}\right)^{N_{Q-}}.$
Figure 2: Leading Lee–Yang zeros $g_{\mathrm{LY}}$ and two-size estimators $\eta_{\mathrm{eff}}(L)$ of the 2+12{+}1D columnar dimerized Heisenberg model at $J_{c}=J_{2}/J_{1}=1.90951$ [54]. (a) $\ln|G|$ along the imaginary-source axis for $L=128$, with red vertical lines marking the leading Lee–Yang zeros. (b) $\eta_{\mathrm{eff}}(L)$ from $(L,2L)$ pairs using the first four zeros. The dashed line shows the $\mathrm{O}(3)$ reference form $\eta+cL^{-\omega}$ with fixed $\eta=0.0375$ and $\omega=0.78$. Inset: leading $g_{\mathrm{LY}}$ versus $1/L$, fitted to $a\,L^{-y_{h}}\!\left(1+bL^{-\omega}\right)$, with fitted $\eta$ defined by $y_{h}=(d+z+2-\eta)/2$. The fit uses fixed $\omega=0.78$ and excludes $L\leq 32$.
Figure 3: Leading Lee–Yang zeros $g_{\mathrm{LY}}$ and two-size estimators $\eta_{\mathrm{eff}}(L)$ of the 2+12{+}1D CBJQ model at $J_{c}=J/Q=0.2174$ in panels (a,b), and at $J_{c}=J/Q=0.2175$ [105] in panels (c,d). $\eta_{\mathrm{eff}}(L)$ is obtained from $(L,2L)$ pairs using the first two zeros. The dashed line shows the first-order reference form $\eta+c_{1}L^{-\omega}+c_{2}L^{-2\omega}$ with fixed $\eta=-1$. Panels (a,c) show zeros from analytic continuation of the Néel field $h$ conjugate to $M^{z}$, and panels (b,d) show zeros from analytic continuation of the plaquette-singlet-solid (PSS) source $\eta$ conjugate to $O^{\mathrm{PSS}}$. Insets: leading $g_{\mathrm{LY}}$ versus $1/L$, fitted to $a\,L^{-y_{h}}$, with fitted $\eta$ defined by $y_{h}=(d+z+2-\eta)/2$. The fits exclude $L\leq 32$.
Figure 4: Leading Lee–Yang zeros $g_{\mathrm{LY}}$ and two-size estimators $\eta_{\mathrm{eff}}(L)$ of the 2+12{+}1D J-Q3 model at $J_{c}=J/Q_{3}=0.67045$ [53, 106, 84] in panels (a,b), and of the 2+12{+}1D J-Q2 model at $J_{c}=J/Q_{2}=0.045$ [72, 81, 78, 66, 84] in panels (c,d). $\eta_{\mathrm{eff}}(L)$ is obtained from $(L,2L)$ pairs using the first two zeros. The dashed line shows the first-order reference form $\eta+c_{1}L^{-\omega}+c_{2}L^{-2\omega}$ with fixed $\eta=-1$. Panels (a,c) show zeros from analytic continuation of the Néel field conjugate to $M^{z}$, and panels (b,d) show zeros from analytic continuation of the columnar VBS source $\xi_{x}$ conjugate to $D_{x}$. Insets: leading $g_{\mathrm{LY}}$ versus $1/L$, fitted to $a\,L^{-y_{h}}$, with fitted $\eta$ defined by $y_{h}=(d+z+2-\eta)/2$. The fits exclude $L\leq 32$.
Our primary targets are the uniform J-Q2 and J-Q3 models, which are canonical DQCP candidates [70, 57, 53, 72, 59, 60, 78, 68, 84, 76], shown in Fig. 1(c) and (d). In these systems, conventional real-axis observables can appear stable over the accessible range of sizes and may yield seemingly converged critical exponents. As we show below, Lee–Yang zeros provide a more direct probe of the finite-size singularity scale and thereby expose slow drift more clearly. The Hamiltonians are
$H_{\text{JQ}_{2}}$ $=-J\sum_{\langle ij\rangle}P_{ij}-Q_{2}\!\!\sum_{\langle ij,kl\rangle\in\square}P_{ij}P_{kl},$ (3)
$H_{\text{JQ}_{3}}$ $=-J\sum_{\langle ij\rangle}P_{ij}-Q_{3}\!\!\sum_{\langle ij,kl,mn\rangle\in\square}P_{ij}P_{kl}P_{mn},$ (4)
where $\langle ij,kl\rangle\in\square$ denotes the two parallel bonds of an elementary plaquette, in either orientation, and $\langle ij,kl,mn\rangle\in\square$ denotes three adjacent parallel bonds on a $2\times 3$ or $3\times 2$ rectangle [70, 53, 78]. Tuning $J/Q_{p}$, with $p\in\{2,3\}$, drives the transition between the Néel AFM and the columnar VBS [70, 57, 53, 72, 59, 78]. To detect VBS order, we probe the bond-supported columnar channels at $\mathbf{Q}=(\pi,0)$ and $(0,\pi)$, using $D_{\mu}=\sum_{\mathbf{r}}(-1)^{r_{\mu}}\,\mathbf{S}_{\mathbf{r}}\!\cdot\!\mathbf{S}_{\mathbf{r}+\hat{\mu}}$. Motivated by the emergent $\mathrm{SO}(5)$ structure discussed for these transitions [59, 60, 93, 78, 68, 84, 76], one may equivalently probe the AFM sector or use plaquette-based definitions of VBS order; here we focus on the bond VBS channel and summarize alternative probes in the Supplemental Material. Implementing the imaginary source term $\mathrm{i}\xi_{\mu}D_{\mu}$, we separate bond insertions of orientation $\mu$ into the two translation-related subclasses distinguished by the sign of $(-1)^{r_{\mu}}$, and denote the corresponding counts by $N_{J+}$ and $N_{J-}$ [16]. [^4] In the SSE representation, $N_{J+}$ and $N_{J-}$ count diagonal $J$-bond operator insertions on bonds of orientation $\mu$, corresponding to the two subclasses with $(-1)^{r_{\mu}}=\pm 1$. The resulting factorized ratio contribution is [16] $R_{J\mu}(\mathrm{i}\xi_{\mu})=\left(1+\mathrm{i}\frac{\xi_{\mu}}{J}\right)^{N_{J+}}\left(1-\mathrm{i}\frac{\xi_{\mu}}{J}\right)^{N_{J-}}.$
At a continuous transition in a given symmetry sector, we denote by $\mathrm{i}g_{\mathbf{Q}}$ the imaginary source coupled to the corresponding order parameter $O_{\mathbf{Q}}$ and $g_{\mathrm{LY}}^{(k)}(L)$ the $k$th Lee–Yang zero in the complex source plane for linear system size $L$. A more detailed discussion of this scaling is given in the Supplemental Material. Since $g_{\mathbf{Q}}$ is the scaling field conjugate to $O_{\mathbf{Q}}$, with renormalization-group eigenvalue $y_{h}$ [23, 65, 11, 32], finite-size scaling gives $g_{\mathrm{LY}}^{(k)}(L)\sim A_{k}\,L^{-y_{h}},$ up to nonuniversal metric factors and irrelevant-operator corrections [32, 37, 34, 38, 17], with nonuniversal amplitude $A_{k}$. Using $y_{h}=(d+z+2-\eta)/2,$ where $d$ is the spatial dimension, $z$ is the dynamical exponent, and $\eta$ is the anomalous dimension of the order parameter, the Lee–Yang zeros therefore provide direct access to $\eta$ [65, 11]. [^5] Throughout this work, $\eta$ extracted from $y_{h}=(d+z+2-\eta)/2$ is used as a finite-size diagnostic. Only in a genuine critical regime does it coincide with the true anomalous dimension; in the coexistence regime, $g_{\mathrm{LY}}(L,\beta)\sim(\beta L^{d})^{-1}$ corresponds to the diagnostic limit $\eta\to-1$. At phase coexistence, by contrast, a small source biases competing phases by an amount proportional to the Euclidean spacetime volume $V_{\tau}=\beta L^{d}$, where $\beta$ is the inverse temperature, since the source term is summed over all sites and imaginary time. The zeros then scale as $g_{\mathrm{LY}}^{(k)}(L,\beta)\sim B_{k}/(\beta L^{d}),$ with nonuniversal amplitude $B_{k}$ [62, 5, 9, 6, 7, 8]. To eliminate $B_{k}$, we use the two-size estimator [^6] Starting from $g_{\mathrm{LY}}^{(k)}(L)\sim A_{k}L^{-y_{h}}$, the ratio for a size pair $(L,2L)$ eliminates $A_{k}$, $g_{\mathrm{LY}}^{(k)}(2L)/g_{\mathrm{LY}}^{(k)}(L)\sim 2^{-y_{h}},$ so $y_{h}=-\ln[g_{\mathrm{LY}}^{(k)}(2L)/g_{\mathrm{LY}}^{(k)}(L)]/\ln 2.$ Combining this with $y_{h}=(d+z+2-\eta)/2$ gives the estimator below. At finite size, corrections to scaling make this a size-dependent effective estimate of $\eta$, rather than its asymptotic value. [54, 17, 92] $\eta_{\mathrm{eff}}^{(k)}(L)=d+z+2+\frac{2}{\ln 2}\ln\left[\frac{g_{\mathrm{LY}}^{(k)}(2L)}{g_{\mathrm{LY}}^{(k)}(L)}\right].$ For $d=2$ and $z=1$, $\eta_{\mathrm{eff}}^{(k)}(L)$ approaches $\eta$ at a continuous transition, whereas for the first-order benchmarks studied here, with $\beta\propto L$, the volume law $g_{\mathrm{LY}}^{(k)}\sim L^{-(d+1)}$ gives the diagnostic value $\eta=-1$ [32, 62, 5, 9, 6, 7]. If the probed channel is not the ordering channel, the free energy remains analytic at $g_{\mathbf{Q}}=0$ in the thermodynamic limit, and the Lee–Yang zeros terminate at a finite edge [25, 11, 26, 4, 64]. In J-Q models, this gives a finite VBS-source (Néel-source) edge in the Néel (VBS) phase, closing only at the transition. We therefore focus on leading-zero scaling near criticality and leave edge behavior and its finite-size crossover to the Supplemental Material [32, 37, 33, 4].
This perspective also clarifies why the leading Lee–Yang zeros can diagnose slow pseudocritical drift more cleanly than conventional real-axis observables in a weakly first-order regime [28, 56]. At first order, coexistence is governed by competition between phases rather than by a single scale-invariant state. Real-axis observables such as order-parameter moments or spin correlations therefore need not approach a unique critical power law and can be strongly affected by phase mixing and analytic backgrounds [62, 5, 9]. By contrast, the leading Lee–Yang zeros directly encode the finite-size singularity scale through their distance from the real axis, and in the coexistence regime they cross over to the characteristic spacetime-volume scaling of a first-order transition. [^7] Near coexistence, finite-volume observables are phase-weighted mixtures of competing states, while thermodynamic response functions, being derivatives of $\ln Z$, are controlled by the nearby Lee–Yang zeros. In the ordering channel this gives $g_{\mathrm{LY}}(L,\beta)\sim(\beta L^{d})^{-1}$. See Supplemental Material for details. The leading zeros can therefore reveal the slow crossover toward coexistence on size windows where conventional diagnostics still appear quasi-critical [17, 92].
Results and Discussion. We first benchmark the Lee–Yang analysis in a system with well-established infrared behavior [36, 71]. The columnar dimerized Heisenberg model hosts a conventional 2+12{+}1D $\mathrm{O}(3)$ critical point [98, 99, 36]. As shown in Fig. 2, the leading zeros in the complex source conjugate to $M^{z}$ are well described by $g_{\mathrm{LY}}(L)=a\,L^{-y_{h}}\!\left(1+bL^{-\omega}\right),$ with $y_{h}=(d+z+2-\eta)/2$ [103, 42, 32, 37, 17]. Fitting with fixed $\omega=0.78$ and excluding $L\leq 32$, we obtain $\eta=0.0368\pm 0.0037$, consistent with the established 3D $\mathrm{O}(3)$ value $\eta=0.0375(5)$ [30, 10, 61]. We also construct $\eta_{\mathrm{eff}}(L)$ from $(L,2L)$ pairs using the first four zeros, which removes the nonuniversal amplitude and exposes finite-size corrections more clearly [54, 17, 92]. The resulting sequence follows the $\mathrm{O}(3)$ reference form $\eta_{\mathrm{eff}}(L)=\eta+cL^{-\omega}$ with fixed $\eta=0.0375$ and $\omega=0.78$ [36, 61]. Notably, $\eta_{\mathrm{eff}}(L)$ drifts upward with increasing $L$, in contrast to the first-order trend, where $g_{\mathrm{LY}}(L)\sim L^{-(d+z)}$ implies $\eta\to-1$ [62, 5, 9, 6, 7].
The CBJQ transition provides a controlled first-order benchmark for our Lee–Yang diagnostics [105, 80, 44]. As shown in Fig. 3, we examine two nearby couplings around the literature estimate $J_{c}/Q=0.2175\pm 0.0001$ [105]. At $J/Q=0.2174$, the leading zeros in both the Néel and PSS channels drift clearly toward the first-order spacetime-volume law $g_{\mathrm{LY}}(L)\sim L^{-(d+z)}$ in the ground-state window $\beta\propto L^{z}$ [62, 5, 9, 6, 7, 8], implying $y_{h}=d+z$ and hence $\eta\to-1$ for $d=2$, $z=1$ [32, 34, 38]. The paired-size flow is likewise well described by the first-order reference form $\eta_{\mathrm{eff}}(L)=\eta+c_{1}L^{-\omega}+c_{2}L^{-2\omega}$ with fixed $\eta=-1$ [62, 5, 9, 6, 7]. At $J/Q=0.2175$, the Néel-channel zeros remain broadly consistent with this trend, while in the PSS channel the first and second zeros split visibly. We interpret this as indicating that $J/Q=0.2175$ lies slightly on the Néel side of the transition, so that a small PSS-channel Lee–Yang edge remains finite on the accessible sizes: the first zero enters the edge-controlled regime earlier than the second, and the corresponding $\eta_{\mathrm{eff}}(L)$ values should therefore not be viewed as following a single critical scaling form [25, 32, 4, 64]. The $J/Q=0.2174$ data thus provide the cleaner first-order benchmark in both channels. Overall, the CBJQ results establish the expected first-order Lee–Yang phenomenology and highlight the sensitivity of Lee–Yang zeros to coexistence versus critical scaling even within a narrow coupling window [105, 80, 62, 6, 7]. This provides a useful first-order baseline for the J-Q cases discussed below [72, 78, 105, 56, 84].
We now turn to the uniform J-Q3 and J-Q2 models at their putative Néel–VBS critical couplings. As shown in Fig. 4, we extract the leading Lee–Yang zeros $g_{\mathrm{LY}}(L)$ from two symmetry-resolved deformations, the complex Néel field $h$ conjugate to $M^{z}$ and the columnar VBS source $\xi_{x}$ conjugate to $D_{x}$ [70, 57, 53, 72, 16]. Although restricted fitting windows admit apparent power laws, the paired-size estimator $\eta_{\mathrm{eff}}(L)$, constructed from the first two zeros, shows a systematic downward drift with increasing $L$ in both channels and in both models, relative to the first-order reference form $\eta_{\mathrm{eff}}(L)=\eta+c_{1}L^{-\omega}+c_{2}L^{-2\omega}$ with fixed $\eta=-1$ [54, 16, 92]. For J-Q3, this behavior is consistent with the weakly first-order or pseudocritical interpretation discussed in the literature, namely a slow flow toward the spacetime-volume law $g_{\mathrm{LY}}(L)\sim L^{-(d+z)}$, equivalently $\eta\to-1$ and $\Delta_{\phi}\to 0$ for $d=2$, $z=1$ [6, 7, 34, 59, 78, 56, 28, 76]. For J-Q2, where the nature of the transition has remained less settled, we find closely analogous Lee–Yang flow in both the Néel and VBS probes, providing direct evidence that J-Q2 exhibits the same slow drift toward the first-order benchmark [72, 81, 78, 66, 84]. The Néel-channel sequence is typically smoother than the VBS-channel sequence in our Lee–Yang analysis, suggesting stronger subleading finite-size effects in the VBS probe. Compared with CBJQ, however, the approach to $\eta=-1$ is much slower in both J-Q2 and J-Q3, consistent with an extended pseudocritical regime in which accessible sizes remain quasi-critical over broad intermediate scales before eventually bending toward coexistence [59, 78, 56, 28, 76, 84]. Overall, the Lee–Yang data support a weakly first-order interpretation of both transitions and show that Lee–Yang zeros provide a particularly sensitive probe of this slow drift, especially in the less settled J-Q2 case.
Conclusion. We introduced a symmetry-resolved Lee–Yang framework for quantum magnets that combines SSE sampling at a symmetry-preserving real anchor Hamiltonian with analytic continuation of conjugate sources to evaluate complex generating functions and extract leading zeros. Benchmark calculations recover conventional $\mathrm{O}(3)$ scaling in the columnar dimerized Heisenberg model and the expected first-order scaling behavior in the CBJQ model. For the uniform J-Q3 and J-Q2 models, we find systematic downward drift of the leading-zero scaling in both Néel and VBS channels. In J-Q3, this behavior is consistent with the weakly first-order or pseudocritical scenario discussed in the literature; in J-Q2, we find closely analogous Lee–Yang flow providing direct evidence for the same slow drift toward the first-order benchmark, with the two-size estimator $\eta_{\mathrm{eff}}$ remaining compatible with a slow approach toward $\eta\to-1$, equivalently $\Delta_{\phi}\to 0$, consistent with an extended pseudocritical regime and a weakly first-order interpretation of the Néel–VBS transition. Our work therefore clarifies a long-standing question from the perspective of complexified partition functions and demonstrates the power of the Lee–Yang zero method.
Acknowledgments. We thank Q. Ye, X. Li, J. Rong and W. Zhu for helpful discussions. Z.W. acknowledges support from the China Postdoctoral Science Foundation under Grant No. 2024M752898. Z.L. is supported by the China Postdoctoral Science Foundation under Grants No. 2024M762935. This project is supported by the Scientific Research Project (No. WU2025B011) and the Start-up Funding of Westlake University. H.Z. is supported by the National Natural Science Foundation of China (Grant No. 12274126). The authors thank the IT service office and the high-performance computing center of Westlake University.